Find the whole quantity if 
  40% of it is 500k km
Find the whole quantity if 
  40% of it is 500k km

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the whole quantity be x, Then,
(40/100) × (x) = 500
X = 500 × (100/40)
X = 500 × (10/4)
X = 500 × 2.5
X = 1250 km

Thanks